Write the equation of the line which has y-intercept (0,5) and is perpendicular to the line with equation y=-3x+1.
Answer by rapaljer(4671)

Remember that the equation of a line is y =mx+b where m is the slope and b is the y intercept.

The slope of the given line, y = -3x+1, is -3.

The slope of a line that is perpendicular to this is 1/3.

Since ths y intercept is 5, b=5, so y=mx+b becomes y = 1/3 x + 5.

R^2 at SCC